Al can paint a room in a hours working alone. Bill can paint a room in b hours working
alone. Working alone, Bill takes two more hours to paint a room than Al. Working together,
they can paint a room in 8 hours. Write down a system of two equations involving the variables
a and b that can be used to find a and b. Solve the system.

Answer by stanbon(48568)

Al can paint a room in "a" hours working alone.
Al rate = 1/a job/hr.
----------------------------
Bill can paint a room in "b" hours working alone.
Bill rate = 1/b job/hr.
----------------------------
Working alone, Bill takes two more hours to paint a room than Al.
b = a + 2
-----------------
Working together, they can paint a room in 8 hours.
Together rate = 1/(a + b) job/hr.
------------------------------------
Write down a system of two equations involving the variables a and b that can be used to find a and b.
EQUATIONS:
b = a + 2
(1/(a+b)) = 1/8
=========================
Cheers,
Stan H.